Use a diagrams to present the following:

1. 1/2 -1/8

2. 3/4 - 1/2

3. 1/2 - 1/8

4. 1/4 - 1/8


1
Expert's answer
2021-07-01T13:39:59-0400

Use a diagrams to present the following:

1. "{1\over2} -{1\over 8}"

Rewrite the question using common denominator of 8

"=({1\over 2}\times{4\over 4})-{1\over 8}"

"={4\over 8}-{1\over 8}"

This can be represented in a diagram as follows



"\therefore" "{4\over 8}-{1\over 8}={3\over 8}"



2."{3\over4} -{1\over 2}"

Rewrite the question using common denominator of 4

"={3\over 4}-({1\over 2}\times{2\over 2})"

"={3\over 4}-{2\over 4}"

This can be represented in a diagram as follows



"\therefore" "{3\over 4}-{2\over 4}={1\over 4}"


3. "{1\over2} -{1\over 8}"

Rewrite the question using common denominator of 8

"=({1\over 2}\times{4\over 4})-{1\over 8}"

"={4\over 8}-{1\over 8}"

This can be represented in a diagram as follows


"\therefore" "{4\over 8}-{1\over 8}={3\over 8}"



4. "{1\over 4 }-{ 1\over 8}"

Rewrite the question using common denominator of 8

"=({1\over 4}\times{2\over 2})-{1\over 8}"

"={2\over 8}-{1\over 8}"

This can be represented in a diagram as follows


"\therefore" "{2\over 8}-{1\over 8}={1\over 8}"